OpenStudy (anonymous):

Use L Hopital's Rule to evaluate the following limits: (I'm having problems with three questions based on the instruction and I'll list them one by one. Really need help pleade ^.^;)

OpenStudy (anonymous):

what do u mean??

OpenStudy (anonymous):

Wait, I'll give the first question out ^.^;

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} 2x/(x+7\sqrt{x})\]

OpenStudy (anonymous):

btw I meant please, not pleade ^.^

OpenStudy (anonymous):

whats the correct answer @lightchaste ?

OpenStudy (anonymous):

@fatinatikah The answer is 0. I got it right by using the conjugate method but I want to know how the steps are supposed to be made by using L Hopital's Rule

OpenStudy (anonymous):

u sure the answer is 0?

OpenStudy (anonymous):

can i know your steps?

OpenStudy (wasiqss):

yes the answer is 0

OpenStudy (anonymous):

can i know the working?

OpenStudy (anonymous):

One hint. Take LCM in the denominator :)

OpenStudy (wasiqss):

differentiating both denominator and nominator gives 2 in numerator and {1+3.5/(x^1/2)}

OpenStudy (anonymous):

yep the answer is 0 Using the conjugate method: \[\lim_{x \rightarrow 0}2x/(x+7\sqrt{x})\]\[\lim_{x \rightarrow 0} 2x(x-7\sqrt{x})/[(x-1\sqrt{x})(x+7\sqrt{x})]\]\[\lim_{x \rightarrow 0}2x(x-7\sqrt{x})/(x ^{2}-49x)\] \[\lim_{x \rightarrow 0} 2(x-7\sqrt{x})/(x-49)\] Replace it with the limit you'll get 0

OpenStudy (wasiqss):

put 0 in equation, in denominator, it becomes \[(1+3.5/0)\]

OpenStudy (wasiqss):

3.5/0 makes it infinity , then infinity +1 =infinity hence the final equation is 2/infinity and anything divided by infinity gives 0

OpenStudy (wasiqss):

did u get it man?

OpenStudy (anonymous):

Doesn't 3.5/0 become indeterminate?

OpenStudy (wasiqss):

indeterminate forms applies to whole equation not a part of it, like in this case

OpenStudy (anonymous):

BTW @fatinatikah, I fumbled on the second line of my equation, it's not \[1\sqrt{x}\] it's \[7\sqrt{x}\] ^.^

OpenStudy (anonymous):

@wasiqss is right : Also it doesn't matter Because 7 * 0 =1 *0=0

OpenStudy (anonymous):

Because 7 * 0 =1 *0=0 -->This is in reference to mistake in question 7 sqrt(x) instead o f 1sqrt(x)

OpenStudy (anonymous):

ahhhh now I understand :)

OpenStudy (anonymous):

@shivam_bhalla owh lol right, but since I'm using conjugate it wouldn't make sense in the working even though the answer is right :)

OpenStudy (anonymous):

Question 2 coming up in 20 mins ay :D

OpenStudy (wasiqss):

i will be gone till then :P

OpenStudy (anonymous):

haha lol sorry, gotta do something urgent at the moment :(

OpenStudy (wasiqss):

aww its fine

OpenStudy (anonymous):

But you know what, I'll just type it in quickly \[\lim_{r \rightarrow 1}a(r ^{n}-1)/(r-1)\] the answer is an, but I got an-1 :(

OpenStudy (anonymous):

Simple. Tell me what is differential of r^n with respect to n.

OpenStudy (anonymous):

Well just for the sake of having a reponse/reply, the third question is: \[\lim_{x \rightarrow \infty} (x -\sqrt{x ^{2}+x})\]

OpenStudy (anonymous):

urm..nr^(n-1)?

OpenStudy (anonymous):

yes. Applying l'hopital rule, (d/dn) ( denominator) = 1 (d/dn) ( a *r^n) = a * (d/dn)( r^n) = a*n*r^(n-1) substitute r=1. So 1^(n-1)=1 a*n is the answer

OpenStudy (anonymous):

owh so is my way incorrect? \[\lim_{r \rightarrow 1} (ar ^{n}-1)/(r-1)\]\[\lim_{r \rightarrow 1}(anr ^{n}-1)/1 \] (after being differentiated) \[an -1\]

OpenStudy (anonymous):

Differentiation of constant i.e -1 in your case is zero :) and Sorry for replying late because I was busy at another problem :)

OpenStudy (anonymous):

lol I typed the first line wrong. it's supposed to be (ar^n -r)

OpenStudy (anonymous):

Well since this is going to be a pretty long post ill just close it >.>